OpenStudy (a1234):

a = 5 and c = 20. Find A and round to two decimal places.

OpenStudy (a1234):

 Is A equal to 14.48 degrees?

OpenStudy (anonymous):

Are you familiar with the law of sines?

OpenStudy (a1234):

Yes, but I don't think they're needed here. sinA = 5/20 A = sin^-1 * 5/20

OpenStudy (anonymous):

It makes things a bit easier `LAW OF SINES:` \[\frac{ \sin(A) }{ a }=\frac{ \sin(B) }{ b }\] Plug in what we know \[\frac{ \sin(90) }{ 20 }=\frac{ \sin(A) }{ 5 }\] \[A=\sin^{-1} (\frac{ 5\sin(90) }{ 20 })\]

OpenStudy (anonymous):

Technically, the law applies to the whole triangle so it should be \[\frac{ \sin(A) }{ a }=\frac{ \sin(B) }{ b }=\frac{ \sin(C) }{ c }\] But you get the idea

OpenStudy (anonymous):

Your answer is correct though.

OpenStudy (anonymous):

Eh, it doesn't make it simpler. But later on you'll see that it's great to have.

OpenStudy (a1234):

Yes, it can be useful, but maybe not in this particular case

OpenStudy (a1234):

Thanks!